a girl is 3 years old and her brother is 3 times older than her. when she turns 7, how old will her brother be?
Answer by drj(1380)

A girl is 3 years old and her brother is 3 times older than her. when she turns 7, how old will her brother be?

Step 1. Let 3*3=9 be age of brother since he is 3 times older than her.

Step 2. Then 7-3=4 years difference between now at age 3 and 4 years from now at age 7.

Step 3. Then her brother will be 9+4=13.

Step 4. ANSWER: In 4 years, her brother is 13 years old.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J